# Feature learning in finite-width Bayesian deep linear networks with multiple outputs and convolutional layers

### v1

@@ Abstract
I Introduction
@@
Significant progress has been made in certain asymptotic regimes, where neural networks enjoy forms of universality that simplify their analysis. In the (lazy-training) infinite-width limit, in particular, universality is manifested in both the training under gradient flow, described by the neural tangent kernel [1], and the Bayesian inference setting [2], where exact relations between neural networks and kernel methods were obtained. However, these results cannot explain the success of modern deep architectures, which perform non-trivial feature selection beyond what is possible in the lazy-training regime [3, 4, 5]. One can circumvent this problem by considering different rescalings of the weights with the width of the layers, using the so-called mean field [6, 7, 8, 9] or the more recent maximal update parametrizations [10].
@@
Despite these insights, neural networks at finite width, which are relevant for applications, display complex and intriguing phenomena that are not captured by infinite-width asymptotics [11, 12, 13, 14, 15, 16]. Addressing these finite-size properties calls for a non-perturbative theory, one which does not rely on large-size limits. Developing such a theory for general neural networks has proven to be quite challenging. However, progress has been made by considering simplified models where the activation functions are linear. Deep linear networks [17, 18, 19] are linear in terms of their inputs but they retain non-linear dependence on the parameters. They provide analytically tractable non-convex problems in parameter space, thus serving as a bridge between idealized infinite-width models and the reality of finite-sized networks.
@@
Here, we address the computation of network statistics in the Bayesian learning setting, equivalent to the statistical mechanics framework employed by physicists [20, 21, 22, 23, 24]. In Bayesian deep learning, the prior over the network’s parameters induces a prior over the network’s outputs (more details are in Sec. II.3). In the (lazy-training) infinite-width limit, and for non-linear activation function, the prior over the outputs is known to be Gaussian both for fully connected architectures [20, 25, 26] and for convolutional architectures [22, 23]. For fully connected networks, precise rates of convergence to normality have been investigated as well; see [27, 28]. The goal of our work is to characterize the non-Gaussian behavior, at finite depth and width, of two classes of deep linear networks: (i) those with fully-connected layers and multiple outputs, and (ii) those with convolutional layers and a single linear readout. Our results, summarized in the next section, provide a way to quantify non-perturbative feature learning effects in these architectures. We will discuss the link between our work and the literature by physicists in Sec. IV.
I.1 Informal statement of the results
@@
In the case of quadratic loss function (Gaussian likelihood), the posterior distribution inherits the properties of the prior. The posterior is again a mixture of Gaussians and this leads to the rather standard equations for the bias and variance of a Gaussian Process [29], with the important difference that they are now random variables (Proposition 3 and 4).
@@
Using the simple parametric dependence of our formulas from the layer widths, we provide an asymptotic analysis in the limit of large width, using the language of Large Deviation Theory. First, we recover the well-known infinite-width limit, showing that the prior is degenerate in this case, since the Wishart ensembles concentrate around the identity matrix (Proposition 5). Second, we consider the so-called feature learning infinite-width limit [3, 6, 30, 31, 10] (which is equivalent to a different re-parametrization of the loss and of the output function) and we precisely show how it provides a way to escape lazy training. The measure over the Wishart ensembles concentrates also in this case, but this time they do so around non-trivial solutions that explicitly depend on the training inputs and labels (Proposition 6).
I.2 Related work
@@
Saxe, McClelland and Ganguli derived exact solutions for the training dynamics of deep linear networks (DLN) in [17, 18]. More recently, Li and Sompolinsky [19] studied deep linear networks in a Bayesian framework, via the equivalent statistical mechanics formulation, and obtained analytical results in the proportional limit, where the common size of the hidden layers $N$ and the size of the training set $P$ are taken to infinity, while keeping their ratio $\alpha=P/N$ fixed. This analysis was later generalized to globally gated DLNs [32], and extended to the data-averaged case, using the replica method from spin glass theory [33].
@@
Hanin and Zlokapa reconsidered the same setting in the case of single-output networks [34], and they derived a non-asymptotic result for the Bayesian model evidence (partition function in statistical mechanics), in terms of Meijer G-functions (the fact that the prior over a single output is related to these special functions was shown in [35]). They employed this result to investigate various asymptotic limits where $N$, $P$, and the depth of the network $L$ are simultaneously taken to infinity at the same rate.
@@
Closely related to the deep linear literature, the authors of Ref. [36] proposed an effective theory for deep non-linear networks in the proportional limit, whose derivation is based on a Gaussian equivalence [37, 38, 39, 40, 41, 42] heuristically justified using Breuer-Major theorems [43, 44, 45]. This effective theory has been tested with great accuracy for one-hidden-layer fully-connected networks with a single output in Ref. [46]. In Refs. [47, 48] the authors reconsidered the same setting in the Bayes optimal and data-averaged case. Other non-perturbative approaches to Bayesian deep non-linear networks include Refs. [49, 50, 51] and the recent renormalization group framework proposed in Refs. [52, 53].
I.3 Kernel renormalization in the proportional limit
@@
Bayesian networks in the infinite-width limit ($N\to\infty$ at finite $P$) are equivalent to Gaussian processes (NNGP), whose covariance is given by a kernel [21]. A powerful view of Bayesian networks in the proportional limit [19, 36, 32] focuses on how the prior is modified layer by layer. For fully-connected architectures with a single output, the NNGP kernel at each layer is rescaled by a real number (a scalar order parameter in statistical mechanics). The order parameters satisfy a set of coupled equations that explicitly depend on the inputs and labels in the training set. This process is referred to as a scalar renormalization of the kernel. In fully-connected networks with multiple outputs, instead, the order parameters are $D\times D$ matrices ($D$ being the size of the readout layer), and they modify the NNGP kernel via a Kronecker product. This process is called kernel shape renormalization [19, 36]. Yet another transformation of the kernel, coined local kernel renormalization, was found for convolutional networks [54], where the NNGP kernel splits into several local components [22], labelled by two additional indices $i,j$ that run across the patches in each element of the training set (e.g., the receptive fields of an image). In the proportional limit, each local component $(i,j)$ of the kernel is modified by an order parameter $Q_{ij}$. An interesting type of kernel shape renormalization was found in globally gated deep linear networks too [32], but the interpretation is different than in the convolutional case presented in [54].
@@
$\mathbf{y}^{\mu}$, $\mathbf{y}_{1:P}$ label (response) of $\mathbf{x}^{\mu}$ in the training set, $\mathbf{y}_{1:P}^{\top}=(\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{P}^{\top})$
@@
$\mathbf{S}^{\mu}$ output of the network in $\mathbb{R}^{D}$, $\mathbf{S}^{\mu}=f_{\theta}(\mathbf{x}^{\mu})$
@@
$S^{\mu}$ output of the network in $\mathbb{R}$, $S^{\mu}=f_{\theta}(\mathbf{x}^{\mu})$
@@ II Problem setting
II.1 Fully-connected deep linear neural networks (FC-DLNs)
@@
$\mathbf{S}_{1:P}=(S^{1}_{1},S^{2}_{1},\dots,S^{P}_{1},\dots,S^{1}_{D},\dots,S^{P}_{D})^{\top}$
the $P$ outputs $(\mathbf{S}^{\mu}=f_{\theta}(\mathbf{x}^{\mu}):\mu=1,\dots,P)$ stacked in a vector. .
II.2 Convolutional deep linear neural networks (C-DLNs)
@@ II.3 Bayesian setting
III Results
III.1 Exact non-asymptotic integral representation for the prior over the outputs
@@
Equivalently, $Q$ has a Wishart distribution with $N$ degrees of freedom and scale matrix $V$ if its law corresponds to the law of $\sum_{i=1}^{N}\mathbf{Z}_{i}\mathbf{Z}_{i}^{\top}$ where $\mathbf{Z}_{i}$ are independent Gaussian vectors in $\mathbb{R}^{D}$ with zero mean and covariance matrix $V$ (see [55] and the references therein). If $D=1$, one has that $\mathcal{S}^{+}_{D}=\mathbb{R}^{+}$ and the Wishart distribution reduces to a Gamma distribution of parameters $\alpha=\frac{N}{2}$, $\beta=\frac{1}{2V}$, that is
[…]
@@
We now show that the prior distribution over $\mathbf{S}_{1:P}$ is a mixture of Gaussians where the covariance matrix is an explicit function of random Wishart matrices. In what follows $A\otimes B$ denotes the Kroneker product of matrices $A$ and $B$. Moreover, we set $\lambda^{*}:=\lambda_{0}\dots\lambda_{L}$.
Proposition 1 (non-asymptotic integral rep. for the prior of FC-DLNs with multiple outputs)
Let $\hat{p}_{\mathrm{prior}}$ be the characteristic function of the outputs $\mathbf{S}_{1:P}$ of a FC-DLN, that is $\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\mathbf{E}[\exp\{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}\}]$. If $\min(N_{\ell}:\ell=1,\dots,L)>D$, then
@@
where $\mathcal{Q}_{L,N}$ is defined by (8), $\mathcal{K}_{\mathrm{FC}}(Q_{1},\dots,Q_{L})=Q^{(L)}(Q_{1},\dots,Q_{L})\otimes(N_{0}\lambda^{*})^{-1}X^{\top}X$, and
@@
In the previous representation one can choose any decomposition of $Q_{\ell}$ in $U_{\ell}^{\top}U_{\ell}$. For convenience in what follows we choose the Cholesky decomposition where $U_{\ell}^{\top}=U_{\ell}^{\top}(Q_{\ell})$ is a lower triangular matrix with positive diagonal entries. This $U_{\ell}^{\top}$ is called Cholesky factor (or Cholesky square root). The Cholesky square root is one to one and continuous from $\mathcal{S}_{D}^{+}$ to its image. When $Q_{\ell}$ has a Wishart distribution the law of it Cholesky square root $U_{\ell}^{\top}$ is the so-called Bartlett distribution (see [56]).
@@
As recalled above, for $D=1$ the $U_{\ell}$’s are independent random variables distributed as the square root of a $\operatorname{Gamma}(N_{\ell}/2,N_{\ell}/2)$ random variable, that is $N_{\ell}{U_{\ell}}^{2}$ is a $\chi^{2}$ with ${N_{\ell}}$ degrees of freedom. In the case of Gaussian likelihood, we can employ this result to obtain an alternative integral representation of the non-asymptotic partition function (Bayesian model evidence) given by Hanin and Zlokapa [34].
While in Ref. [34] the partition function is given in terms of Meijer G-functions, here it is expressed as an $L-$dimensional integral over a set of order parameters $Q_{\ell}$. Using our alternative representation, it is possible to compare and find an equivalence with the results (in the proportional limit $P,N_{\ell}\to\infty$ at fixed $\alpha_{\ell}=P/N_{\ell}$) found in Refs. [19, 36] (see also Discussion in Section IV). Note that the link between Meijer G-functions and the integral representation based on Gamma functions was first established by Zavatone-Veth and Pehlevan in Ref. [35] in the limited setting of individual training patterns (i.e. for the prior over a single input vector). It should be noted that when $L=1$, Proposition 1 shows that $\mathbf{S}_{1:P}$ is a mixture of Gaussians with a Wishart matrix as covariance. It is known that the resulting distribution is the so-called multivariate Student’s $t$-distribution (see, e.g., Representation B in [57]). One can show that the characteristic function of a multivariate Student’s $t$ can be expressed in term of Macdonald functions [58]. Since the Macdonald function is recovered as a special case of a Meijer G-function, this clarify (at least for $L=1$) the link between our stochastic representation and the Fourier representation given in [35] for $P=1$.
@@
Proposition 2 (non-asymptotic integral rep. for the prior of C-DLNs with single output)
For a C-DLN with $\min(C_{\ell}:\ell=1,\dots,L)>N_{0}$, the characteristic function $\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\mathbf{E}[\exp\{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}\}]$ of the outputs $\mathbf{S}_{1:P}$ is given by
@@
Note that the map $\mathcal{T}$ acts as on a four indices tensor, providing a partial trace wrt to indices $r,\,s$. This tensor is known in the literature as local covariance matrix from the seminal work on the Bayesian infinite-width limit in convolutional networks [22]. The same partial trace structure has been also identified in the proportional limit for non-linear one-hidden-layer networks in [54], where the authors coined the term local kernel renormalization to indicate the re-weighting of the infinite-width local NNGP kernel found in [22].
III.2 Posterior distribution and predictor’s statistics in the case of squared error loss function

### v2

@@ Abstract
Keywords: Deep learning theory, Bayesian deep linear networks, Convolutional layers, Feature learning, Gaussian mixtures
1 Introduction
@@
Significant progress has been made in certain asymptotic regimes, where neural networks enjoy forms of universality that simplify their analysis. In the (lazy-training) infinite-width limit, in particular, universality is manifested in both the training under gradient flow, described by the neural tangent kernel (Jacot et al., 2018), and the Bayesian inference setting (Lee et al., 2018b), where exact relations between neural networks and kernel methods were obtained. However, these results cannot explain the success of modern deep architectures, which perform non-trivial feature selection beyond what is possible in the lazy-training regime (Chizat et al., 2019; Lewkowycz et al., 2021; Brown et al., 2020). One can circumvent this problem by considering different rescalings of the weights with the width of the layers, using the so-called mean field (Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2022; Sirignano and Spiliopoulos, 2020; Chizat and Bach, 2018) or the more recent maximal update parametrizations (Yang and Hu, 2021).
@@
Despite these insights, neural networks at finite width, which are relevant for applications, display complex and intriguing phenomena that are not captured by infinite-width asymptotics (Ciceri et al., 2024; Ingrosso and Goldt, 2022; Petrini et al., 2023b, a; Sclocchi and Wyart, 2024; Sclocchi et al., 2023). Addressing these finite-size properties calls for a non-perturbative theory, one which does not rely on large-size limits. Developing such a theory for general neural networks has proven to be quite challenging. However, progress has been made by considering simplified models where the activation functions are linear. Deep linear networks (Saxe et al., 2014, 2019; Li and Sompolinsky, 2021) are linear in terms of their inputs but they retain non-linear dependence on the parameters. They provide analytically tractable non-convex problems in parameter space, thus serving as a bridge between idealized infinite-width models and the reality of finite-sized networks.
@@
Here, we address the computation of network statistics in the Bayesian learning setting, equivalent to the statistical mechanics framework employed by physicists (Neal, 1996; Lee et al., 2018a; Novak et al., 2019; Garriga-Alonso et al., 2019; Engel and Van den Broeck, 2001). In Bayesian deep learning, the prior over the network’s parameters induces a prior over the network’s outputs (more details are in Sec. 2.3). In the (lazy-training) infinite-width limit, and for non-linear activation function, the prior over the outputs is known to be Gaussian both for fully connected architectures (Neal, 1996; de G. Matthews et al., 2018; Hanin, 2023) and for convolutional architectures (Novak et al., 2019; Garriga-Alonso et al., 2019). For fully connected networks, precise rates of convergence to normality have been investigated as well (see Favaro et al., 2023; Trevisan, 2023). The goal of our work is to characterize the non-Gaussian behavior, at finite depth and width, of two classes of deep linear networks: (i) those with fully-connected layers and multiple outputs, and (ii) those with convolutional layers and a single linear readout. Our results, summarized in the next section, provide a way to quantify non-perturbative feature learning effects in these architectures. We will discuss the link between our work and the literature by physicists in Sec. 4.
1.1 Informal statement of the results
@@
In the case of quadratic loss function (Gaussian likelihood), the posterior distribution inherits the properties of the prior. The posterior is again a mixture of Gaussians and this leads to the rather standard equations for the bias and variance of a Gaussian Process (Rasmussen and Williams, 2005), with the important difference that they are now random variables (Proposition 6 and 7).
@@
Using the simple parametric dependence of our formulas from the layer widths, we provide an asymptotic analysis in the limit of large width, using the language of Large Deviation Theory. First, we recover the well-known infinite-width limit, showing that the prior is degenerate in this case, since the Wishart ensembles concentrate around the identity matrix (Proposition 8). Second, we consider the so-called feature learning infinite-width limit (Chizat et al., 2019; Mei et al., 2018; Geiger et al., 2021, 2020; Yang and Hu, 2021), which is equivalent to a different re-parametrization of the loss and of the output function, and we precisely show how it provides a way to escape lazy training. The measure over the Wishart ensembles concentrates also in this case, but this time they do so around non-trivial solutions that explicitly depend on the training inputs and labels (Proposition 10).
1.2 Related work
@@
Saxe et al. (2014, 2019) derived exact solutions for the training dynamics of deep linear networks (DLN). More recently, Li and Sompolinsky (2021) studied deep linear networks in a Bayesian framework, via the equivalent statistical mechanics formulation, and obtained analytical results in the proportional limit, where the common size of the hidden layers $N$ and the size of the training set $P$ are taken to infinity, while keeping their ratio $\alpha=P/N$ fixed. This analysis was later generalized to globally gated DLNs (Li and Sompolinsky, 2022), and extended to the data-averaged case, using the replica method from spin glass theory (Zavatone-Veth et al., 2022).
@@
Hanin and Zlokapa (2023) reconsidered the same setting in the case of single-output networks, and they derived a non-asymptotic result for the Bayesian model evidence (partition function in statistical mechanics), in terms of Meijer G-functions (the fact that the prior over a single output is related to these special functions was shown in Zavatone-Veth and Pehlevan, 2021). They employed this result to investigate various asymptotic limits where $N$, $P$, and the depth of the network $L$ are simultaneously taken to infinity at the same rate.
@@
Closely related to the deep linear literature, Pacelli et al. (2023) proposed an effective theory for deep non-linear networks in the proportional limit, whose derivation is based on a Gaussian equivalence (Goldt et al., 2020, 2022; Gerace et al., 2020; Hu and Lu, 2023; Gerace et al., 2024; Aguirre-López et al., 2024) heuristically justified using Breuer-Major theorems (Breuer and Major, 1983; Bardet and Surgailis, 2013; Nourdin et al., 2011). This effective theory has been tested with great accuracy for one-hidden-layer fully-connected networks with a single output in Baglioni et al., 2024. Camilli et al., 2023 and Cui et al., 2023 reconsidered the same setting in the Bayes optimal and data-averaged case. Other non-perturbative approaches to Bayesian deep non-linear networks include Naveh and Ringel (2021); Seroussi et al. (2023); Fischer et al. (2024) and the recent renormalization group framework proposed in Howard et al. (2024a, b).
1.3 Kernel renormalization in the proportional limit
@@
Bayesian networks in the infinite-width limit ($N\to\infty$ at finite $P$) are equivalent to Gaussian processes (NNGP), whose covariance is given by a kernel (Lee et al., 2018a). A powerful view of Bayesian networks in the proportional limit (Li and Sompolinsky, 2021; Pacelli et al., 2023; Li and Sompolinsky, 2022) focuses on how the prior is modified layer by layer. For fully-connected architectures with a single output, the NNGP kernel at each layer is rescaled by a real number (a scalar order parameter in statistical mechanics). The order parameters satisfy a set of coupled equations that explicitly depend on the inputs and labels in the training set. This process is referred to as a scalar renormalization of the kernel. In fully-connected networks with multiple outputs, instead, the order parameters are $D\times D$ matrices ($D$ being the size of the readout layer), and they modify the NNGP kernel via a Kronecker product. This process is called kernel shape renormalization (Li and Sompolinsky, 2021; Pacelli et al., 2023). Yet another transformation of the kernel, coined local kernel renormalization, was found for convolutional networks in Aiudi et al. (2023), where the NNGP kernel splits into several local components (Novak et al., 2019), labelled by two additional indices $i,j$ that run across the patches in each element of the training set (e.g., the receptive fields of an image). In the proportional limit, each local component $(i,j)$ of the kernel is modified by an order parameter $Q_{ij}$. An interesting type of kernel shape renormalization was found in globally gated deep linear networks too (Li and Sompolinsky, 2022), but the interpretation is different than in the convolutional case presented in Aiudi et al. (2023).
@@
$\mathbf{y}^{\mu}$, $\mathbf{y}_{1:P}$ label (response) of $\mathbf{x}^{\mu}$ in the training set, $\mathbf{y}_{1:P}^{\top}=(\mathbf{y}^{1\top},\dots,\mathbf{y}^{P\top})$
@@
$\mathbf{S}^{\mu}$, $\mathbf{S}_{1:P}$ output of the network in $\mathbb{R}^{D}$, $\mathbf{S}^{\mu}=f_{\theta}(\mathbf{x}^{\mu})$, $\mathbf{S}_{1:P}^{\top}=(\mathbf{S}^{1\top},\dots,\mathbf{S}^{P\top})$
@@
$S^{\mu}$, $\mathbf{S}_{1:P}$ output of the network in $\mathbb{R}$, $S^{\mu}=f_{\theta}(\mathbf{x}^{\mu})$, $\mathbf{S}_{1:P}^{\top}=(S^{1},\dots,S^{P})$
@@ 2 Problem setting
2.1 Fully-connected deep linear neural networks (FC-DLNs)
@@
$\mathbf{S}_{1:P}=(S^{1}_{1},\dots,S^{1}_{D},S^{2}_{1},\dots,S^{P}_{1},\dots,S^{P}_{D})^{\top}$
the $P$ outputs $(\mathbf{S}^{\mu}=f_{\theta}(\mathbf{x}^{\mu}):\mu=1,\dots,P)$ stacked in a vector.
2.2 Convolutional deep linear neural networks (C-DLNs)
@@ 2.3 Bayesian setting
3 Results
Our main results are collected in this section.
3.1 Exact non-asymptotic integral representation for the prior over the outputs
@@
Equivalently, $Q$ has a Wishart distribution with $N$ degrees of freedom and scale matrix $V$ if its law corresponds to the law of $\sum_{i=1}^{N}\mathbf{Z}_{i}\mathbf{Z}_{i}^{\top}$ where $\mathbf{Z}_{i}$ are independent Gaussian vectors in $\mathbb{R}^{D}$ with zero mean and covariance matrix $V$ (see Eaton, 2007 and the references therein). If $D=1$, one has that $\mathcal{S}^{+}_{D}=\mathbb{R}^{+}$ and the Wishart distribution reduces to a Gamma distribution of parameters $\alpha=\frac{N}{2}$, $\beta=\frac{1}{2V}$, that is
$\mathcal{W}_{1}(Q|V,N)=\frac{1}{(2V)^{N/2}\Gamma(N/2)}Q^{\frac{N}{2}-1}e^{-\frac{1}{2V}Q}\quad\text{for $Q>0$}.$
@@
We now show that the prior distribution over $\mathbf{S}_{1:P}$ is a mixture of Gaussians where the covariance matrix is an explicit function of random Wishart matrices. In what follows $A\otimes B$ denotes the Kronecker product of matrices $A$ and $B$. Moreover, we set $\lambda^{*}:=\lambda_{0}\dots\lambda_{L}$.
Proposition 1
(non-asymptotic integral representation for the prior of FC-DLNs with multiple outputs) Let $\hat{p}_{\mathrm{prior}}$ be the characteristic function of the outputs $\mathbf{S}_{1:P}$ of a FC-DLN, that is $\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\mathbf{E}[\exp\{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}\}]$. If $\min(N_{\ell}:\ell=1,\dots,L)>D$, then
@@
where $\mathcal{Q}_{L,N}$ is defined by (8), $\mathcal{K}_{\mathrm{FC}}(Q_{1},\dots,Q_{L})=(N_{0}\lambda^{*})^{-1}X^{\top}X\otimes Q^{(L)}(Q_{1},\dots,Q_{L})$, and
@@
In the previous representation one can choose any decomposition of $Q_{\ell}$ in $U_{\ell}^{\top}U_{\ell}$. For convenience in what follows we choose the Cholesky decomposition where $U_{\ell}^{\top}=U_{\ell}^{\top}(Q_{\ell})$ is a lower triangular matrix with positive diagonal entries. This $U_{\ell}^{\top}$ is called Cholesky factor (or Cholesky square root). The Cholesky square root is one to one and continuous from $\mathcal{S}_{D}^{+}$ to its image. When $Q_{\ell}$ has a Wishart distribution the law of it Cholesky square root $U_{\ell}^{\top}$ is the so-called Bartlett distribution (see Kshirsagar, 1959).
@@
As recalled above, for $D=1$ the $U_{\ell}$’s are independent random variables distributed as the square root of a $\operatorname{Gamma}(N_{\ell}/2,N_{\ell}/2)$ random variable, that is $N_{\ell}{U_{\ell}}^{2}$ is a $\chi^{2}$ with ${N_{\ell}}$ degrees of freedom. In the case of Gaussian likelihood, we can employ this result to obtain an alternative integral representation of the non-asymptotic partition function (Bayesian model evidence) given by Hanin and Zlokapa (2023).
While in Hanin and Zlokapa (2023) the partition function is given in terms of Meijer G-functions, here it is expressed as an $L-$dimensional integral over a set of order parameters $Q_{\ell}$. Using our alternative representation, it is possible to compare and find an equivalence with the results (in the proportional limit $P,N_{\ell}\to\infty$ at fixed $\alpha_{\ell}=P/N_{\ell}$) found in Li and Sompolinsky (2021) and Pacelli et al. (2023) (see also Discussion in Section 4). Note that the link between Meijer G-functions and the integral representation based on Gamma functions was first established in Zavatone-Veth and Pehlevan (2021) in the limited setting of individual training patterns (i.e. for the prior over a single input vector). It should be noted that when $L=1$, Proposition 1 shows that $\mathbf{S}_{1:P}$ is a mixture of Gaussians with a Wishart matrix as covariance. It is known that the resulting distribution is the so-called multivariate Student’s $t$-distribution (see, e.g., Representation B in Lin, 1972). One can show that the characteristic function of a multivariate Student’s $t$ can be expressed in term of Macdonald functions (Joarder and Alam, 1995). Since the Macdonald function is recovered as a special case of a Meijer G-function, this clarify (at least for $L=1$) the link between our stochastic representation and the Fourier representation given in Zavatone-Veth and Pehlevan (2021) for $P=1$.
@@
Proposition 5
(non-asymptotic integral representation for the prior of C-DLNs with single output) For a C-DLN with $\min(C_{\ell}:\ell=1,\dots,L)>N_{0}$, the characteristic function $\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\mathbf{E}[\exp\{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}\}]$ of the outputs $\mathbf{S}_{1:P}$ is given by
@@
Note that the map $\mathcal{T}$ acts as on a four indices tensor, providing a partial trace wrt to indices $r,\,s$. This tensor is known in the literature as local covariance matrix from the seminal work Novak et al. (2019) on the Bayesian infinite-width limit in convolutional networks. The same partial trace structure has been also identified in the proportional limit for non-linear one-hidden-layer networks in Aiudi et al. (2023), where the authors coined the term local kernel renormalization to indicate the re-weighting of the infinite-width local NNGP kernel found in Novak et al. (2019).
3.2 Posterior distribution and predictor’s statistics in the case of squared error loss function